Call Us demand function is Q = 5,000 – 0.5P and the MC = $100. Given that TFC is 20% of TR at the revenue maximizing level

a. At present Call Us is operating where its MR = 0. Explain the implication.

b. Derive the AFC (average fix cost) at the profit maximizing level.
Answer #1

a. At the present Call US is operating where MR=0 that means Total Revenue o Call Us is maximum. They are operating at the optimal profit maximising level. If they produce little more of the product thus MR will be negative (downward sloped MR) then TR will start to fall.

b. Demand Function Q=5000-0.5P

where Q=output

P= price per unit

Total Revenue (TR)= Quantity* Price per unit

= 5000 - 0.5P * P

= 5000P - 0.5P2

Marginal Revenue (MR) = d(TR/P) = 5000-P

Given, MR=0, P=5000 and  Q= 5000- 0.5 *5000 =2500

TR=( 5000 - 0.5*5000) * 5000 = 1250000

TFC = 20% of TR = (20/100)*1250000 = 250000

AFC = TFC/ Q = 250000/2500 = 100